### v2

# smICA: an open source repository for mapping the concentration of fluorescently labeled molecules in living cells on the basis of confocal imaging combined with time-correlated single-photon counting

Abstract
Advanced microscopy techniques are essential for visualizing and tracking cellular components and molecules in biomedical research. However, conventional fluorescence microscopy methods often struggle with accurately measuring molecule concentrations in cells. To overcome these limitations, we introduce a novel approach that integrates laser scanning confocal microscopy with time-correlated single photon counting (TCSPC), supported by an open-source analysis tool called smICA (single-molecule Image to Concentration Analyzer). Our method, validated against traditional fluorescence correlation spectroscopy (FCS), offers enhanced accuracy in determining fluorescent molecule concentrations, particularly in cases where molecules are immobile or unevenly distributed. This is demonstrated using fluorescently labeled mRNA in living cells, highlighting the approach’s effectiveness.
Introduction
Advanced microscopes are necessary for quantitative biological studies, especially when visualizing
cells, localizing and tracking single molecules, or analyzing their concentration. Fluorescence microscopy,
particularly confocal microscopy, has become a standard tool in biomedical research. Numerous techniques have
evolved from traditional laser scanning confocal microscopy (LSCM), including fluorescence correlation
spectroscopy
(FCS), 11, 4, 3, 10 raster image correlation spectroscopy
(RICS),1 and fluorescence lifetime imaging microscopy (FLIM). These techniques are based on
time-correlated single photon counting (TCSPC),
which enables the tracking of photons emitted by fluorescence molecules, one by one, in time.
There are various ways to monitor the concentration of fluorescent molecules. Qualitatively,
fluorophore’s concentration in the given region of the sample can be deduced from the fluorescence intensity as it is proportional to concentration.
However, obtaining quantitative measurements is more challenging, and FCS is the more appropriate technique of
choice. The disadvantage of the correlation-based methods is that they are limited to mobile molecules and are
blind to molecules that are, for some reason, immobile. Moreover, the traditional FCS measurements only offer
information from a single spot. Althought, there were attempts to preform multipoint
FCS,14, 6, 18 the technique reqires customized equipment and is difficult to attempt on the
commercailly available setups. Therefore, the representative FCS measurements can only be conducted under the
assumption of homogenous analyte distribution within the sample. The limitations, including the
restriction to
mobile species, spatial constraints, and compartmentalization make quantifying concentrations in living cells particularly
challenging. This difficulty arises because target biomolecules often localize within specific subcellular
structures and organelles, resulting in an uneven distribution throughout the cell. Analyzing a map or the
probe’s concentration distribution is more desirable in such cases.
An example of a qualitative approach to studying the number of fluorescent molecules registered on a given
image was proposed by Vukojevic et al..16 The authors used Poissonian statistics
to measure the average number of molecules in the observation volume without temporal or spatial correlation
of the images. Due to strong deviations from the Possonian statistics lying on its ground, the method is
limited to two orders of magnitude of the analyte concentration (1-100 nM). The method described by Vukojevic
et al.16 has the same grounds as the Number and Brightness (N&B) method
proposed by the group of Gratton2 using the relation between the mean and variance of
the fluorescence signal to the molecular brightness and number of fluorescent molecules. The advantage of the
N&B is that the molecular brightness of the molecule of interest does not need to be known as apriori as it
is determined from the fluorescence fluctuation signal. Moreover, the method discriminates between the mobile
and immobile molecule fractions. Both methods, however, do not consider factors such as the autofluorescence
of biological samples or the cross-talk between fluorescent dyes
when measurements are performed in the multi-color mode.
We propose an alternative aproach for determination of the concntration of fluorophores instead of
their number. We combine the laser scanning confocal microscope
with the time-correlated single photons counting (TCSPC) acquisition method. We assume that the
molecular brightness of the analyte is known
apriori and is spatially and temporarily constant. We provide an open-source smICA (single-molecule Image to Concentration
[…]
https://github.com/TKmist/smICA. The bundle consists of three scripts:
EXTRACT_AND_FILTER_PTU, REWRITE_ROI, and Phot2Conc. See
https://github.com/TKmist/smICA/blob/main/README.md
for more details. and perform validation test comparing the
concentration determined with the proposed methodology with the well known FCS method. Finaly we provide the
proof of
concept results for the method by applying it to
study the concentration of fluorescently labeled mRNA in living cells. We observed that many fluorescently
labeled mRNA molecules are entrapped in subcellular structures or organelles, making them immobile and
significantly non-uniformly distributed. While we do not discuss the biological aspects of this phenomenon, it
helps us to show that our imaging method outdoes the FCS measurements, which may lead to significant
underestimation of the mean mRNA concentration.
Materials and methods
Cell culture
Human lung carcinoma cells (A549 line) were obtained from the American Type Culture Collection (ATCC). The cells were cultured in DMEM (Institute of Immunology and Experimental Technology, Wrocław, Poland) supplemented with 10% fetal bovine serum, L-glutamine (2 mM), penicillin (100 g/ml), and streptomycin (100 g/ml) from Sigma-Aldrich. The culture was maintained at 37°C in a 5% humidified atmosphere. Passaging was carried out at approximately 80% confluence every 2-3 days using 0.25% Trypsin-EDTA solution (Sigma-Aldrich) and PBS (Sigma-Aldrich).
mRNA injections
The A549 cells for microinjections were plated 1 or 2 days before the measurements on a 35 mm glass bottom cell culture dish (glass coverslip bottom, 0.17 µm, ibidi, Germany) to grow to approximately 50-70 % confluence. Microinjections were performed by the Femtojet system (Eppendorf) using disposable sterile injection glass capillaries (Femtotip II, Eppendorf, 0.5 µm inner and 0.7 µm outer diameter). The injection time was set to 0.1 s. Before FCS measurements but after microinjections, the growth medium in the culture dish was washed out and replaced with PBS buffer containing Ca2+ and Mg2+ ions (Sigma-Aldrich).
Experimental setups
The measurements were performed using the NIKON C1 and NIKON A1 confocal microscopes. Both setups are additionally equipped with the TCSPC acquisition unit (LSM Upgrade-kit from PicoQuant, Germany). The LSM units are composed of three pulsed diode lasers (485, 561, and 641 nm of wavelength) controlled by the PDL 828 "SEPIA II" controller, the TCSPC module PicoHarp 300, and the pair of single photon avalanche diode - SPAD, detectors with appropriate filters. We used the PLAN APO water immersion objective x60 NA 1.20 in all experiments. Both confocal systems are equipped with an incubator, allowing for temperature control. Measurements in buffers were performed at 298 K, while experiments were conducted at 309 K.
Materials
The fluorescently labeled mRNA (Cy5-m7GRNAegfp, stained with sulfo-Cy5 at the 5’ Cap) was synthesized and purified according to the previously published procedure12. Briefly, mRNA was in vitro transcribed in the presence of azido-modified cap analog (N3-m7GpppG) and subsequently isolated with spin columns. After labeling with sulfo-Cy5-DBCO, the product was isolated with HPLC and characterized with gel electrophoresis. The TRITC labeled dextran of molecular weights 4.4 kDa, 40 kDa and 155 kDa were bought from Sigma-Aldrich.
Fluorescence correlation spectroscopy - callibration, brighntess and cocnentration
Before each experiment, calibration measurements were performed using the fluorescence correlation
spectroscopy technique (FCS). Depending on the laser and the fluorophore under study, we performed calibration using
Rhodamine 110 (for 485 nm laser line), Rhodamine B (561 nm laser line) and Alexa Fluor 647 (for 641 nm
laser line). The diffusion coefficients for
both dyes at 25 degrees were taken from the reference 8. For measurements in living cells, the
calibration measurements were performed at 36 degrees, and the diffusion coefficients were recalculated,
including the change in temperature and viscosity. We get the structure parameter (fixed in target
FCS measurements), the width of the focal volume, , and focal volume, from the calibration
measurements.
Then we performed FCS measurements of our samples to determine the molecular brightness, of used probes. We obtained
the average number of molecules in the confocal focus, from analysing FCS auto-correlation curve and photon-counts per
second,
from the raw FCS time-trace signal. To determine the
for a given FCS measurement, we calculated
the cumulative sum of detected photons for each second of measurement and fitted the result with homogeneous linear function (). is equal to the slope of the fitted function. The brightness is given as . For each sample we
performed a series of FCS measurements and calculated an average value of from the series.
FCS measuremewnts in cells were performed in the cell cytoplasm trying to avoid the endoplasmic reticulum and the
bright spots where the accumulation of mRNA molecules is expected. The stipped region of the cell marked in Figure
1 depicts region of the cell where typical FCS measurements were conducted.
Concnetration imaging methodology
The single-photon fluorescence imaging data, if not properly analyzed, consists of artifacts
originating from the photophysics (triplet states or photobleaching), samples (dye crosstalk,
autofluorescence), and detectors (afterpulsing). Quantitative analysis of dye concentration requires
recognition of each artifact source and reduction of their influence on the experimental data. Below, we
describe our strategy to tackle the challenge.
The scheme of the methodology is depicted in Figure 2.
First, we use standard FCS measurements to calibrate the confocal volume and determine the molecular
brightness, , of the fluorophore used in further experiments. The molecular brightness depends on the power
of the incident light and the experimental setup settings. Therefore, measuring under the same conditions as
the final experiments is recommended. We further assume that in the sample, does not change, neither
spatially nor temporally.
After calibration, we perform the LSCM imaging combined with the SPAD detectors utilizing the TSCPC
method. Standard LSCM hardware provides filtration via optical filters and dichroic mirrors. The TCSPC method
used in the time-tagged time-resolved (TTTR) mode,17 additionally, allows filtering fluorescence
data by methods
relying on the fluorescence decay pattern such as fluorescence lifetime filtering (commonly used in
fluorescence lifetime correlation spectroscopy)5, 9 or the pulsed
interleaved excitation combined with the time gating.
The raw TTTR data are extracted from the binary .ptu files, optionally filtered, integrated over all
registered frames, and combined into the image. Each image pixel contains a total number of photons registered
at a given spot in the sample. Next, for each pixel, we calculate the effective number of
detected molecules,, using the fluorophore’s molecular brightness, . The mean
number of molecule detected at a given spot is given as , where stands for the pixel dwell time and is the number of frames acquired.
Finally, knowing the dimensions of the effective focal volume, we calculate the mean concentration per pixel;
, where
is the Avogadro number and corresponds to the volume of the focal spot obtained from the of FCS
calibration measurements.
The graphical user interface
The tool for concentration imaging consists of three scripts. Although the scripts perform simple tasks, they
are designed to be as intuitive as possible. We aim to use them by users who are not trained in programming
and allow them to analyze large amounts of data files in a repeatable manner. Here, we briefly describe the
[…]
scripts are based on the Dear PyGUI - the graphical user interface toolkit(GUI) for Python, making them fast,
dynamic, and intuitive.
The first GUI script, EXTRACT_AND_FILTER_PTU, is dedicated to extracting the raw TTTR data from the
binary .ptu file where the data acquired by the commercial, FLIM/FCS dedicated SymPhoTime software are
stored. The data are extracted with the help of the readPTU_FLIM library
[…]
desgned to read the
.ptu files. Read TTTR data gathered in different channels are split. In each channel, the user can
filter the data using the time gating method or by applying the statistical filters according to the algorithm
described in references 5 and 9. As an output,
the script returns a series of files consisting of the following: (i) an array the number of photons
registered in each pixel of the unfiltered image (one file per channel), (ii) an array for intensity
values (number of photons after filtration) per pixel (one file per channel), (iii) an array for
fluorescence average lifetime values per pixel after filtration (one file per channel), (iv) the
decay pattern for each channel (separate files), (v) a .png file that can be further used to set the
region of interest - ROI.
The second GUI script is called REWRITE_ROI. The use of this script is optional and is related to
the ROI. The ROI used in the further analysis can be selected directly in the SymPhoTime at the data
acquisition stage. However, when a lot of data is acquired and time between subsequent data points is
crucial, selecting the ROI after the data acquisition is more desired. In our experiments, we used the ImageJ
software to read the .png files generated by the EXTRACT_AND_FILTER_PTU script and then create the ROI as a
mask, which was further saved as the text image. If the ROI data are not created using the SymPhoTime
software, using the REWRITE_ROI script is necessary. Our script transcribes the ImageJ-generated files into
the SymPhoTime format.
The third script performs all calculations required to translate the number of photons into the
concentration. The user can select the set of roi files (by choosing the folder) corresponding to the analyzed
ptu files. The FCS calibration parameters, including the focal volume structure parameter, ,
the lateral size of the focal volume, , and the molecular brightness, , can be manually added or
loaded from the .json file. The calculations are performed for all files selected by the user. As an
output, the script returns the text files containing (i) an array of the number of molecules per
pixel, (ii) an array of the mean concentration of molecules per pixel, (iii) the table in
the user-desired format (.csv, .xlsx, .dat, .pickle111This is
the binary file storing the pandas DataFrame.). The output table consisting of the File name,
channel number, the mean number of counts (photons), mean number of molecules per pixel, mean concentration
per pixel, and the corresponding standard deviations. All mean values correspond to the average over the
entire image or ROI (if selected).
Results and discussion
Validation and tests
To validate the determination of the concentration,
we designed experiments where we changed the
fluorophore’s concentration, power of the laser (measured at the sample), the pixel dwell time - ,
and the size of the pixel or, more precisely, the scan velocity defined as
.222The variable controlled experimentally in the hardware setup is . We
assume, however, that the effective size of the pixel is equal to the focal volume iluminated
within the time . The focal volume is determined from the FCS calibration measurements.
We measured a well-defined probe - the dextran polymer of molecular weight of 155 kg/mol labeled with
the TRITC dye. This polymer is characterized by the hydrodynamic radius of the order of 9 nm. The
measurements
on dextrans were performed in
water at 25℃. The concentration of the polymers varied from 6 to 100 nM. The laser
power, , was
changed between 5 to 40 W, which corresponded to the peak focal irradiance ,333, where stands for the radius of the focal volume. varying
from 0.07 to
0.5 mW/m2. The pixel dwell time was changed between 3.12 s to 60 s, and the scan velocity
changed between 0.4 nm/s to 64.1 nm/s. At low polymer concentrations, laser power, and short ,
the signal-to-noise ratio, SNR, was close to unity and the probability of detection of a photon from
the fluorescent molecule was below unity. For example, the average number of photons, =0.013, expected
to be observed per one dextran molecule at =5 W (=4260 photons/molecule/s) and = 3.12
s. For the highest laser power of =40 W (=12300 photons/molecule/s) and = 60 s,
=0.74. To increase the chance of detecting any photon, we integrated the signal over ten frames in
post-acquisition processing, increasing the expected values by ten times.
Post-acquisition processing also includes the filtration of unwanted photons originating from
afterpulsing, the sample’s autofluorescence, or the dye-dye crosstalk. Technically, the filtration procedure
is performed on the raw TCSPC signal before the over-all-frames integration.
In the EXTRACT_AND_FILTER_PTU, GUI, we implemented two methods of photons filtering, both of which relied
on TCSPC data. The first method is based on the fluorescence lifetime decay
pattern 5, 9, where the properly calculated weights are applied
to each registered photon. The filtering was applied to the FITC-labeled dextran data and compared to the
non-filtered data; see Figure
3a. We calculated the concentration of fluorescent dextran molecules for
each image pixel and then averaged over all pixels. We further compared the mean value to the concentration
obtained from imaging, with the one etermined using FCS, on the same sample at the same
concentration and laser power conditions, by calculating their ratio, . We found that independently on the experimental
conditions (laser’s power, pixel dwell time, pixel size or concentration) the oscilates around unity. The distribution of
the values for all measurements are depicted as violin plots in Figure 3b.
During image analysis, we also
checked the influence of the number of pixels used in analyis (size of the region of interest - ROI). We
found that the filtration procedure caused over-filtration of the fluorescence signal, manifested as
systematically lowered values of the ratio. Interestingly the size of the ROI do not influence the
results significantly. We used several sizes of square regions of interest differing in the number of
pixels, , varying from 162 to 2562 pixels (=2562 refers to the full frame). In both
filtered and non-filtered data, the ROI size did not influence the mean value of the ratio, but data
variance decreases with the ROI’s dimension. For the ROI size , the differences in
data distribution were unrecognizable; cf. Figure 3b. The mean value of
the ratio for all non-filtered data is close to 1.05, with the square root of the variance ()
at the level of 0.23 for the smallest ROI and not higher than 0.19 for . On the
contrary, the filtered data for all ROI sizes exhibits values not higher than 0.17, but the mean
value of the ratio did not exceed 0.78.
The second type of filter implemented in the EXTRACT_AND_FILTER_PTU GUI is based on the time-gating method, which is dedicated to cases when the crosstalk between fluorophores occurs. To verify the photon filtering efficiency under the dye-dye crosstalk conditions, we performed test measurements on the cells where two types of fluorophores were present, the EGFP and the TRITC-labeled dextran. The EGFP was synthenized in cells after inection of the mRNA encoding the EGFP sequence. The TRITC-labeled dextran was mixed with with the non-labeled mRNA at known proportins and injected into cells. Mixing of the mRNA and dextran allowed us to estimate the concentration of the inected mRNA. To separate the excitation events between the laser lines, we operated in the pulse interleaved excitation, PIE, mode, allowing us to shift the consecutive laser’s pulses in time. We used the shift between the equal to 25 ns, and the delay time between pulses of the same laser of 50 ns. We analyzed the data from sixteen different cells. The power of the lasers was set to 10 W for the green laser ( = 561 nm) and 2 W for the blue laser ( = 485 nm). The images were acquired for = 399 nm, = 8 s and integrated over 100 frames.
Figures 4a and b show exemplary fluorescence decay patterns of the TRITC-labeled dextran and the EGFP, acquired with two SPAD detectors corresponding to the TRITC and EGFP channels. The pulse of the green laser is shifted by +25 ns with respect to the pulse of the blue laser. To filter out the photons originating from the dye-dye crosstalk recorded in the TRITC channel, we selected the decay-time range from 25-50 ns; Figure 4a. Consequently, the range of decay times for the EGFP channel was chosen as between 0 and 25 ns; 4b. Figure 4c compares the for the filtered and non-filtered signals for FITC-labeled dextran and eGFP, respectively. Although the differences in the ratio for the filtered and non-filtered are not very pronounced, the drop in the values agrees with the differences in intensities from the crosstalking fluorophores.
Example of implementation
We apply the described methodology to the quantitative measurements of the concentration of the fluorescently labeled mRNA and its changes upon the time after injection. The cells were injected with CAP-modified mRNA fluorescently labeled with the sulfo-Cy5 dye.12 The mRNA was encoding the green fluorescent protein. The initial concentration of injected mRNA was around 100 nM and varied from cell to cell. Concentration imaging was performed on three cells. Between the image acquisition, we performed FCS measurements to determine the concentration of mRNA molecules freely diffusing along the cytoplasm.
[…]
Conclusions
We presented a methodology and the GUI scripts for mapping the fluorophore’s concentration in living cells. The methodology requires the calibration of the confocal volume utilizing standard FCS measurement and assumes that the brightness of the fluorophore is constant. Meeting both requirements allows us to determine the mean concentration per pixel, averaged over the data acquisition time. Therefore, an output contains information about the average concentration over the entire cell/ROI and the concentration map across the cell/ROI. As a proof of concept, we applied the method to monitor the changes in the concentration of fluorescently labeled mRNA in living cells. Observed changes in mRNA suggest that the presented methodology is a promising tool for quantitative single-cell studies. Finally, narrowing ROI to specific organelles can provide insight into more detailed topics, including but not limited to protein expression, degradation of biomolecules, monitoring of enzymatic reactions, and investigation of probes’ affinity of probes for specific cellular targets.
Acknowledgments
Research funded by the Polish Science Fund within the framework of the Virtual Research Institute; grant WIB-1/2020-O11 - WIB_HERO.
Authors contribution
TK: Methodology, Software, Validation, Formal analysis, Investigation, Data Curation, Visualization, Supervision, Writing - Original Draft, Writing - Review & Editing. GB: Investigation, Data Curation, Methodology, Validation, Writing - Original Draft, Writing - Review & Editing. JM: Investigation, Visualization, Data Curation, Writing - Review & Editing. KK: Investigation, Data Curation, Methodology, Software (tesitng), Writing - Original Draft, Writing - Review & Editing. MP: Methodology, Software (tesitng), Writing - Review & Editing. AM: Resources. JJ: Resources, Funding acquisition. RH: Funding acquisition, Writing - Review & Editing.